# Cutoff for Contingency Table Random Walks

### v1

@@ 6.2 Bounding Decay in Characteristic Function for Large $\theta$
In this section, we bound the decay of the characteristic function $\Phi_{Y}(\theta)$ of the $n\times n$ contingency table random walk in the regime $\|\theta\|_{\infty}>\dfrac{1}{\sqrt{q}}$. We will eventually plug in $\theta=\dfrac{2\pi y}{q}$, $y\in({\mathbb{Z}}\cap[-q/2,q/2))^{(n-1)^{2}}$. First, note that we can write the characteristic function $\Phi_{Y}(\theta)$ as
$\frac{1}{{\binom{n}{2}}^{2}}\left(\sum_{\begin{subarray}{c}i>\ell\\ j>k\end{subarray}}\cos(\theta_{ij}-\theta_{ik}-\theta_{\ell j}+\theta_{\ell k})+\sum_{\begin{subarray}{c}i>\ell\\ j\end{subarray}}\cos(\theta_{ij}-\theta_{\ell j})+\sum_{\begin{subarray}{c}j>k\\ i\end{subarray}}\cos(\theta_{ij}-\theta_{ik})+\sum_{i,j}\cos(\theta_{ij})\right).$
We may assume that $\theta\in[-\pi,\pi]^{(n-1)^{2}}$.
@@
Suppose that $\dfrac{1}{\sqrt{q}}<\|\theta\|_{\infty}\leq\dfrac{\pi}{4}$. Recall that $Y_{1}$ has covariance matrix $\dfrac{4}{n^{2}}\Gamma$. Since $\cos(x)\leq 1-\dfrac{2}{25}x^{2}$ for $x\in[-\pi,\pi]$, by the second-order Taylor expansion of $\Phi_{Y}$,
$\Phi_{Y}(\theta)\leq 1-\frac{2}{25}\theta^{T}\left(\frac{4}{n^{2}}\Gamma\right)\theta=1-\frac{8}{25n^{2}}\theta^{T}\Gamma\theta.$ (52)
Next, note that $\Gamma$ has eigenvalue $\dfrac{1}{(n-1)^{2}}$ corresponding to the eigenvector $\mathbbm{1}_{(n-1)^{2}}$, eigenvalue $\dfrac{n}{(n-1)^{2}}$ with an eigenspace equal to $\operatorname{span}(\{\delta_{i.}\}_{i=2}^{n-1}\cup\{\delta_{.j}\}_{j=2}^{n-1})$, where for $i,j\in\{2,\ldots,n-1\}$ and $\ell,k\in\{1,\dots,n-1\}$, we define
$\delta_{i.}(\ell,k):=\begin{cases}-1&\ell=1\\ 1&\ell=i\\ 0&\ell\notin\{1,i\}\end{cases},\hskip 25.0pt\delta_{.j}(\ell,k):=\begin{cases}-1&k=1\\ 1&k=j\\ 0&k\notin\{1,j\}\end{cases}.$
Additionally, $\Gamma$ has an eigenvalue $\left(1+\dfrac{1}{n-1}\right)^{2}$ with an eigenspace equal to
$\operatorname{span}(\{\delta_{ij}\}|i,j\in\{2,\ldots,n-1\}),$
where we define the vectors
$\delta_{ij}(\ell,k):=\begin{cases}-1&\left(\ell=1\text{ and }k=1\right)\text{ or }\left(\ell=i\text{ and }k=j\right)\\ 1&\left(\ell=1\text{ and }k=j\right)\text{ or }\left(\ell=i\text{ and }k=1\right)\\ 0&\ell\notin\{1,i\}\text{ or }k\notin\{1,j\}\end{cases}.$
@@
[…]
If $|\overline{\theta}|\geq\dfrac{1}{8\sqrt{q}}$, then we immediately have
$\theta^{T}\Gamma\theta\geq\frac{1}{(n-1)^{2}}\|\overline{\theta}\mathbbm{1}_{(n-1)^{2}}\|_{2}^{2}=\frac{1}{64q}.$ (53)
So we assume that $|\overline{\theta}|<\dfrac{1}{8\sqrt{q}}$. Also, suppose that there exists either a column or row (WLOG we will assume a row $i$) such that $|\overline{\theta}_{i.}|\geq\dfrac{1}{4\sqrt{q}}$. Then there must be some additional row $i^{\prime}$ such that
@@
Since $w:=\delta_{i.}-\delta_{i^{\prime}.}$ is an eigenvector for $\Gamma$ with eigenvalue $\dfrac{n}{(n-1)^{2}}$, we know that the projection of $\theta$ onto the span of $w$ is equal to
@@
$=\left(\frac{\overline{\theta}_{i.}-\overline{\theta}_{i^{\prime}.}}{2}\right)^{2}\frac{n}{(n-1)^{2}}\|w\|_{2}^{2}$
$\geq\frac{1}{4}\left(\frac{1}{8\sqrt{q}}\right)^{2}\frac{2n}{n-1}$
@@
$|\overline{\theta}|<\frac{1}{8\sqrt{q}},\hskip 15.0pt|\overline{\theta}_{i.}|<\frac{1}{4\sqrt{q}}\text{ for }i\in\{1,\ldots,n\},\quad\text{and}\quad|\overline{\theta}_{.j}|<\frac{1}{4\sqrt{q}}\text{ for }j\in\{1,\ldots,n\}.$
Let $\hat{\theta}$ be the projection of $\theta$ onto $\operatorname{span}\left(\mathbbm{1}_{(n-1)^{2}},\{\delta_{i.}\}_{i=1}^{\infty},\{\delta_{.j}\}_{j=1}^{\infty}\right)$. It follows that
@@
Now with (52), (53), (55), and (59), we conclude that when $\dfrac{1}{\sqrt{q}}<\|\theta\|_{\infty}\leq\dfrac{\pi}{4}$,
$\Phi_{Y}(\theta)\leq 1-\frac{8}{25n^{2}}\frac{1}{128q}=1-\frac{1}{400n^{2}q}.$ (60)
@@
We will now bound the characteristic function $\Phi_{Y}(\theta)$ for the regime where $\|\theta\|_{\infty}\in[\frac{\pi}{4},\pi]$. Recall that we can write $\Phi_{Y}(\theta)$ as
$\frac{1}{{\binom{n}{2}}^{2}}\left(\sum_{\begin{subarray}{c}i>\ell\\ j>k\end{subarray}}\cos(\theta_{ij}-\theta_{ik}-\theta_{\ell j}+\theta_{\ell k})+\sum_{\begin{subarray}{c}i>\ell\\ j\end{subarray}}\cos(\theta_{ij}-\theta_{\ell j})+\sum_{\begin{subarray}{c}j>k\\ i\end{subarray}}\cos(\theta_{ij}-\theta_{ik})+\sum_{i,j}\cos(\theta_{ij})\right)$ (61)
Let
$\mathcal{A}:=\{(i,j)\in\{1,\ldots,n-1\}^{2}:|\theta_{ij}|>1/8\}.$
Suppose that $|{\mathcal{A}}|\geq\dfrac{(n-1)^{2}}{32}$. Then we can write
$\Phi_{Y}(\theta)$ $\leq 1-\dfrac{\sum_{(i,j)\in\mathcal{A}}1-\cos(\theta_{ij})}{{\binom{n}{2}}^{2}}$ (62)
$\leq 1-\dfrac{(n-1)^{2}(1-\cos(1/8))}{32{\binom{n}{2}}^{2}}$
$=1-\frac{1-\cos(1/8)}{8n^{2}},$
@@
Next, suppose that $|{\mathcal{A}}|\leq\dfrac{(n-1)^{2}}{32}$. Pick $(i^{*},j^{*})$ such that $|\theta_{i^{*}j^{*}}|=\|\theta\|_{\infty}$. Define
:=${\mathcal{B}}:=\{i\in\{1,2,\dots,n-1\}:|\theta_{ij^{*}}|>1/4\}$
@@
${\mathcal{C}}:=\{j\in\{1,2,\dots,n-1\}:|\theta_{i^{*}j}|>1/4\}.$
Suppose $|{\mathcal{B}}|\leq\dfrac{n-1}{2}$ and $|{\mathcal{C}}|\leq\dfrac{n-1}{2}$. We look at the terms in (61) of the form $\cos(\theta_{ij}-\theta_{ik}-\theta_{\ell j}+\theta_{\ell k})$. We would replace all the terms by $1$ except those of the form $\cos(\theta_{i^{*}j^{*}}-\theta_{i^{*}j}-\theta_{ij^{*}}+\theta_{ij})$. There are at least $\dfrac{(n-1)^{2}}{4}$ terms of this form with $i\in{\mathcal{B}}^{c}$ and $j\in{\mathcal{C}}^{c}$. Since $|{\mathcal{A}}|\leq\dfrac{(n-1)^{2}}{32}$, there are at least $\dfrac{7(n-1)^{2}}{32}$ of terms $\cos(\theta_{i^{*}j^{*}}-\theta_{i^{*}j}-\theta_{ij^{*}}+\theta_{ij})$ such that $i\in{\mathcal{B}}^{c},j\in{\mathcal{C}}^{c}$, and $|\theta_{ij}|\leq 1/8$. Therefore,
$\Phi_{Y}(\theta)$ $\leq 1-\frac{\sum 1-\cos(\theta_{i^{*}j^{*}}-\theta_{i^{*}j}-\theta_{ij^{*}}+\theta_{ij})}{\binom{n}{2}^{2}}$ (63)
$\leq 1-\frac{\frac{7}{32}(n-1)^{2}(1-\cos(\pi/4-5/8))}{\binom{n}{2}^{2}}$
$=1-\frac{7(1-\cos(\pi/4-5/8))}{8n^{2}}$
where on the first line, the summation ranges over all the pairs $(i,j)$ such that $i\in{\mathcal{B}}^{c},j\in{\mathcal{C}}^{c}$, and $|\theta_{ij}|\leq 1/8$. Now, we suppose that $|{\mathcal{A}}|\leq\dfrac{(n-1)^{2}}{32}$ and $|{\mathcal{C}}|>\dfrac{n-1}{2}$. Note that the case of $|{\mathcal{A}}|\leq\dfrac{(n-1)^{2}}{32}$ and $|{\mathcal{B}}|>\dfrac{n-1}{2}$ would follow in an analogous way. We look at the terms in (61) of the form $\cos(\theta_{i^{*}j}-\theta_{\ell j})$ where $j\in{\mathcal{C}}$ and $\ell\neq i^{*}$. There are at least $\dfrac{(n-1)(n-2)}{2}\geq\dfrac{(n-1)^{2}}{4}$ pairs of $(\ell,j)$ such that $j\in{\mathcal{C}}$ and $\ell\neq i^{*}$. So, there are at least $\dfrac{7(n-1)^{2}}{32}$ pairs of $(\ell,j)$ such that $j\in{\mathcal{C}}$, $\ell\neq i^{*}$, and $|\theta_{\ell j}|\leq 1/8$. Therefore,
$\Phi_{Y}(\theta)$ $\leq 1-\frac{\sum 1-\cos(\theta_{i^{*}j}-\theta_{\ell j})}{\binom{n}{2}^{2}}$ (64)
$\leq 1-\frac{\frac{7}{32}(n-1)^{2}(1-\cos(1/8))}{\binom{n}{2}^{2}}$
$=1-\frac{7(1-\cos(1/8))}{8n^{2}}$
@@ 6.2.3 Combining the Two Regimes of Large $\theta$
Now combining (60), (62), (63), and (64), we conclude that for all large $n$ and $q$,
$\Phi_{Y}(\theta)\leq 1-\frac{1}{400n^{2}q}$ (65)
when $\theta\in[-\pi,\pi]^{(n-1)^{2}}$ satisfies $\|\theta\|_{\infty}>\dfrac{1}{\sqrt{q}}$. So for $y\in({\mathbb{Z}}\cap[-q/2,q/2))^{(n-1)^{2}}$ with $\|y\|_{\infty}>\dfrac{\sqrt{q}}{2\pi}$,
$\exp\left(2t\left(\Phi_{Y}\left(\frac{2\pi y}{q}\right)-1\right)\right)\leq\exp\left(-\frac{t}{200n^{2}q}\right).$ (66)
By (50) with $r=4$, $\sigma^{2}=\dfrac{4(n-1)^{2}}{n^{2}}$, the mixing time is
$t=\dfrac{n^{2}q^{2}\log(\sqrt{\alpha}(n-1))}{8\pi^{2}}\left(1-\dfrac{4}{3q}\right)^{-1},$ (67)
@@
$\exp\left(2t\left(\Phi_{Y}\left(\frac{2\pi y}{q}\right)-1\right)\right)\leq(n-1)^{-\dfrac{q}{1600\pi^{2}}}$ (68)
when $\|y\|_{\infty}\geq\dfrac{\sqrt{q}}{2\pi}$. Similar to (30), the number of elements $y\in({\mathbb{Z}}\cap[-q/2,q/2))^{(n-1)^{2}}$ with $\|y\|_{\infty}>\dfrac{\sqrt{q}}{2\pi}$ is ${\mathcal{O}}(q^{(n-1)^{2}})$. Thus if $q\gg n^{2}$, then at the mixing time (67), we conclude that as $n\to\infty$,

### v2

@@ 6.2 Bounding Decay in Characteristic Function for Large $\theta$
In this subsection, we bound the decay of the characteristic function $\Phi_{Y}(\theta)$ of the $m\times n$ contingency table random walk in the regime $\|\theta\|_{\infty}>\dfrac{1}{\sqrt{q}}$. We will eventually plug in $\theta=\dfrac{2\pi y}{q}$, $y\in({\mathbb{Z}}\cap[-q/2,q/2))^{(m-1)(n-1)}$. First, note that we can write $\binom{m}{2}\binom{n}{2}\Phi_{Y}(\theta)$ as
$\sum_{\begin{subarray}{c}i>\ell\\ j>k\end{subarray}}\cos(\theta_{ij}-\theta_{ik}-\theta_{\ell j}+\theta_{\ell k})+\sum_{\begin{subarray}{c}i>\ell\\ j\end{subarray}}\cos(\theta_{ij}-\theta_{\ell j})+\sum_{\begin{subarray}{c}j>k\\ i\end{subarray}}\cos(\theta_{ij}-\theta_{ik})+\sum_{i,j}\cos(\theta_{ij}).$
We may assume that $\theta\in[-\pi,\pi]^{(m-1)(n-1)}$.
@@
Suppose that $\dfrac{1}{\sqrt{q}}<\|\theta\|_{\infty}\leq\dfrac{\pi}{4}$. Recall that $Y$ has covariance matrix $\dfrac{4}{mn}\Gamma$. Since $\cos(x)\leq 1-\dfrac{2}{25}x^{2}$ for $x\in[-\pi,\pi]$, by the second-order Taylor expansion of $\Phi_{Y}$,
$\Phi_{Y}(\theta)\leq 1-\frac{2}{25}\theta^{T}\left(\frac{4}{mn}\Gamma\right)\theta=1-\frac{8}{25mn}\theta^{T}\Gamma\theta\leq 1-\frac{8}{25\kappa_{2}n^{2}}\theta^{T}\Gamma\theta$ (57)
Recall from Section 6.1 that $\Gamma$ has eigenvalue $\dfrac{1}{(m-1)(n-1)}$ corresponding to the eigenvector $\mathbbm{1}_{(m-1)(n-1)}$, eigenvalue $\dfrac{m}{(m-1)(n-1)}$ with an eigenspace equal to $\operatorname{span}(\{\delta_{i.}\}_{i=2}^{m-1})$, and eigenvalue $\dfrac{n}{(m-1)(n-1)}$ with an eigenspace $\operatorname{span}(\{\delta_{.j}\}_{j=2}^{n-1}))$. Additionally, $\Gamma$ has an eigenvalue $\dfrac{mn}{(m-1)(n-1)}$ with an eigenspace equal to
$\operatorname{span}(\{\delta_{ij}\}|i\in\{2,\ldots,m-1\},j\in\{2,\ldots,n-1\}).$
@@
$\overline{\theta}:=\frac{\sum_{i,j}\theta_{ij}}{(m-1)(n-1)},\hskip 15.0pt\overline{\theta}_{i.}:=\frac{\sum_{j=1}^{n-1}\theta_{ij}}{n-1},\hskip 15.0pt\overline{\theta}_{.j}:=\frac{\sum_{i=1}^{m-1}\theta_{ij}}{m-1}.$
If $|\overline{\theta}|\geq\dfrac{1}{8\sqrt{q}}$, then
$\theta^{T}\Gamma\theta\geq\frac{1}{(m-1)(n-1)}\|\overline{\theta}\mathbbm{1}_{(m-1)(n-1)}\|_{2}^{2}=\frac{1}{64q}.$ (58)
So we assume that $|\overline{\theta}|<\dfrac{1}{8\sqrt{q}}$. Also, suppose that there exists either a column or row (we will first assume a row $i$) such that $|\overline{\theta}_{i.}|\geq\dfrac{1}{4\sqrt{q}}$. Then there must be some additional row $i^{\prime}$ such that
@@
Since $w:=\delta_{i.}-\delta_{i^{\prime}.}$ is an eigenvector for $\Gamma$ with eigenvalue $\dfrac{m}{(m-1)(n-1)}$, we know that the projection of $\theta$ onto the span of $w$ is equal to
@@
$=\left(\frac{\overline{\theta}_{i.}-\overline{\theta}_{i^{\prime}.}}{2}\right)^{2}\frac{m}{(m-1)(n-1)}\|w\|_{2}^{2}$
$\geq\frac{1}{4}\left(\frac{1}{8\sqrt{q}}\right)^{2}\frac{2m}{m-1}$
@@
Now if there is a column $j$ such that $|\overline{\theta}_{.j}|\geq\dfrac{1}{4\sqrt{q}}$, then a similar argument gives
$\theta^{T}\Gamma\theta\geq\frac{1}{128q}.$ (61)
@@
$|\overline{\theta}|<\frac{1}{8\sqrt{q}},\hskip 15.0pt|\overline{\theta}_{i.}|<\frac{1}{4\sqrt{q}}\text{ for }i\in\{1,\ldots,m\},\quad\text{and}\quad|\overline{\theta}_{.j}|<\frac{1}{4\sqrt{q}}\text{ for }j\in\{1,\ldots,n\}.$
Let $\hat{\theta}$ be the projection of $\theta$ onto $\operatorname{span}\left(\mathbbm{1}_{(m-1)(n-1)},\{\delta_{i.}\}_{i=1}^{\infty},\{\delta_{.j}\}_{j=1}^{\infty}\right)$. It follows that
@@
Now with (57), (58), (60), (61), and (65), we conclude that when $\dfrac{1}{\sqrt{q}}<\|\theta\|_{\infty}\leq\dfrac{\pi}{4}$,
$\Phi_{Y}(\theta)\leq 1-\frac{8}{25\kappa_{2}n^{2}}\frac{1}{128q}=1-\frac{1}{400\kappa_{2}n^{2}q}.$ (66)
@@
We will now bound the characteristic function $\Phi_{Y}(\theta)$ for the regime where $\|\theta\|_{\infty}\in[\frac{\pi}{4},\pi]$. Recall that we can write $\binom{m}{2}\binom{n}{2}\Phi_{Y}(\theta)$ as
$\sum_{\begin{subarray}{c}i>\ell\\ j>k\end{subarray}}\cos(\theta_{ij}-\theta_{ik}-\theta_{\ell j}+\theta_{\ell k})+\sum_{\begin{subarray}{c}i>\ell\\ j\end{subarray}}\cos(\theta_{ij}-\theta_{\ell j})+\sum_{\begin{subarray}{c}j>k\\ i\end{subarray}}\cos(\theta_{ij}-\theta_{ik})+\sum_{i,j}\cos(\theta_{ij})$ (67)
Let $n]:=\{1,\dots,n\}$ and
$\mathcal{A}:=\{(i,j)\in[m-1]\times[n-1]:|\theta_{ij}|>1/8\}.$
Suppose that $|{\mathcal{A}}|\geq\dfrac{(m-1)(n-1)}{32}$. Then we can write
$\Phi_{Y}(\theta)$ $\leq 1-\dfrac{\sum_{(i,j)\in\mathcal{A}}1-\cos(\theta_{ij})}{{\binom{m}{2}\binom{n}{2}}}$ (68)
$\leq 1-\dfrac{(m-1)(n-1)(1-\cos(1/8))}{32{\binom{m}{2}\binom{n}{2}}}$
$=1-\frac{1-\cos(1/8)}{8mn},$
@@
Next, suppose that $|{\mathcal{A}}|\leq\dfrac{(m-1)(n-1)}{32}$. Pick $(i^{*},j^{*})$ such that $|\theta_{i^{*}j^{*}}|=\|\theta\|_{\infty}$. Define
${\mathcal{B}}:=\{i\in[m-1]:|\theta_{ij^{*}}|>1/4\}$
@@
${\mathcal{C}}:=\{j\in[n-1]:|\theta_{i^{*}j}|>1/4\}.$
Suppose $|{\mathcal{B}}|\leq\dfrac{m-1}{2}$ and $|{\mathcal{C}}|\leq\dfrac{n-1}{2}$. We look at the terms in (67) of the form $\cos(\theta_{ij}-\theta_{ik}-\theta_{\ell j}+\theta_{\ell k})$. We would replace all the terms by $1$ except those of the form $\cos(\theta_{i^{*}j^{*}}-\theta_{i^{*}j}-\theta_{ij^{*}}+\theta_{ij})$. There are at least $\dfrac{(m-1)(n-1)}{4}$ terms of this form with $i\in{\mathcal{B}}^{c}$ and $j\in{\mathcal{C}}^{c}$. Since $|{\mathcal{A}}|\leq\dfrac{(m-1)(n-1)}{32}$, there are at least $\dfrac{7(m-1)(n-1)}{32}$ of terms $\cos(\theta_{i^{*}j^{*}}-\theta_{i^{*}j}-\theta_{ij^{*}}+\theta_{ij})$ such that $i\in{\mathcal{B}}^{c},j\in{\mathcal{C}}^{c}$, and $|\theta_{ij}|\leq 1/8$. Therefore,
$\Phi_{Y}(\theta)$ $\leq 1-\frac{\sum 1-\cos(\theta_{i^{*}j^{*}}-\theta_{i^{*}j}-\theta_{ij^{*}}+\theta_{ij})}{\binom{m}{2}\binom{n}{2}}$ (69)
$\leq 1-\frac{\frac{7}{32}(m-1)(n-1)(1-\cos(\pi/4-5/8))}{\binom{m}{2}\binom{n}{2}}$
$=1-\frac{7(1-\cos(\pi/4-5/8))}{8mn}$
where on the first line, the summation ranges over all the pairs $(i,j)$ such that $i\in{\mathcal{B}}^{c},j\in{\mathcal{C}}^{c}$, and $|\theta_{ij}|\leq 1/8$. Now, we suppose that $|{\mathcal{A}}|\leq\dfrac{(m-1)(n-1)}{32}$ and $|{\mathcal{C}}|>\dfrac{n-1}{2}$. Note that the case of $|{\mathcal{A}}|\leq\dfrac{(m-1)(n-1)}{32}$ and $|{\mathcal{B}}|>\dfrac{m-1}{2}$ would follow in an analogous way. We look at the terms in (67) of the form $\cos(\theta_{i^{*}j}-\theta_{\ell j})$ where $j\in{\mathcal{C}}$ and $\ell\neq i^{*}$. There are at least $\dfrac{(n-1)(m-2)}{2}\geq\dfrac{(m-1)(n-1)}{4}$ pairs of $(\ell,j)$ such that $j\in{\mathcal{C}}$ and $\ell\neq i^{*}$. So, there are at least $\dfrac{7(m-1)(n-1)}{32}$ pairs of $(\ell,j)$ such that $j\in{\mathcal{C}}$, $\ell\neq i^{*}$, and $|\theta_{\ell j}|\leq 1/8$. Therefore,
$\Phi_{Y}(\theta)$ $\leq 1-\frac{\sum 1-\cos(\theta_{i^{*}j}-\theta_{\ell j})}{\binom{m}{2}\binom{n}{2}}$ (70)
$\leq 1-\frac{\frac{7}{32}(m-1)(n-1)(1-\cos(1/8))}{\binom{m}{2}\binom{n}{2}}$
$=1-\frac{7(1-\cos(1/8))}{8mn}$
@@ 6.2.3 Combining the Two Regimes of Large $\theta$
Now since $m\leq\kappa_{2}n$, combining (66), (68), (69), and (70), we conclude that for all large $n$ and $q$,
$\Phi_{Y}(\theta)\leq 1-\frac{1}{400\kappa_{2}n^{2}q}$ (71)
when $\theta\in[-\pi,\pi]^{(m-1)(n-1)}$ satisfies $\|\theta\|_{\infty}>\dfrac{1}{\sqrt{q}}$. So for $y\in({\mathbb{Z}}\cap[-q/2,q/2))^{(m-1)(n-1)}$ with $\|y\|_{\infty}>\dfrac{\sqrt{q}}{2\pi}$,
$\exp\left(2t\left(\Phi_{Y}\left(\frac{2\pi y}{q}\right)-1\right)\right)\leq\exp\left(-\frac{t}{200\kappa_{2}n^{2}q}\right).$ (72)
By (54) with $r=4$, $\sigma^{2}=\dfrac{4(m-1)(n-1)}{mn}$, the mixing time is
$t=\dfrac{mnq^{2}\log(\alpha(m-1)(n-1))}{16\pi^{2}}\left(1-\dfrac{4}{3q}\right)^{-1},$ (73)
@@
$\exp\left(2t\left(\Phi_{Y}\left(\frac{2\pi y}{q}\right)-1\right)\right)\leq\left((m-1)(n-1)\right)^{-\dfrac{\kappa_{1}q}{3200\kappa_{2}\pi^{2}}}$ (74)
when $\|y\|_{\infty}\geq\dfrac{\sqrt{q}}{2\pi}$. Similar to (34), the number of $y\in({\mathbb{Z}}\cap[-q/2,q/2))^{(m-1)(n-1)}$ with $\|y\|_{\infty}>\dfrac{\sqrt{q}}{2\pi}$ is ${\mathcal{O}}(q^{(m-1)(n-1)})$. Thus if $q\gg mn$, then at the mixing time (73), we conclude that as $n\to\infty$,